OpenStudy (yoyo2443):

What is the equation in standard form of a perpendicular line that passes through (5,1)? The line perpendicular points (5,1) is y=-4x+4

OpenStudy (lord_box):

What's true about the slopes of perpendicular lines?

OpenStudy (yoyo2443):

That their slopes are negative reciprocals...?

OpenStudy (lord_box):

So what would be the negative reciprocal of the slope of -4x + 4?

OpenStudy (yoyo2443):

\[\frac{ 1 }{ 4 }\]

OpenStudy (lord_box):

Right, so you know the slope of the function, and a point on the function. You can construct an equation using this information.

OpenStudy (yoyo2443):

I'm not sure how I would apply it for the standard form though because some information is missing (standard form is Ax + By= C x and y being the points and the rest.. I'm not so sure..)

OpenStudy (lord_box):

First, put it in point slope form, and we'll work it from there.

OpenStudy (yoyo2443):

alright so.. \[y-1=\frac{ 1 }{ 4 }(x-5)\]

OpenStudy (lord_box):

Good, now take it to y = mx + b form.

OpenStudy (lord_box):

By first distributing the 1/4 and moving the -1 over you can accomplish this.

OpenStudy (yoyo2443):

\[y-1=\frac{ 1 }{ 4 }x -1 \frac{ 1 }{ 4 }\] something like that?

OpenStudy (lord_box):

Yes, move the -1 to the right side

OpenStudy (yoyo2443):

\[y= \frac{ 1 }{ ?4}x - \frac{ 1 }{ ?4}\] ?

OpenStudy (yoyo2443):

... forgot to remove the question marks there... My bad

OpenStudy (lord_box):

Yeah, now just move the 1/4x to the left side and it should look familiar.

OpenStudy (yoyo2443):

\[x \frac{ 1 }{ 4 }-y=\frac{ 1 }{ 4 }\]

OpenStudy (lord_box):

Looks good.

OpenStudy (yoyo2443):

Awesome, thank you!

OpenStudy (lord_box):

You're welcome.